### v1

# Wigner Cat Phases: A finely tunable system for exploring the transition to quantum chaos

Abstract
The transition to chaos for quantum dynamics is quantified via a finely tunable mixed random matrix ensemble. The mixed Gaussian Orthogonal Ensemble (mGOE) forms a pedagogically accessible family of systems in simulating Many-Body Localization (MBL) transitions. It can be tuned from chaotic to localized and heavy-tailed localized phases in a continuous fashion, providing an opportunity to explore new phases. We numerically study how the spectral properties of mGOE evolve during these transitions. Characterization of transition to quantum chaos is computed and analyzed via empirical spectral density, nearest-neighbor spacing, and adjacent gap ratios with statistical uncertainty quantifications that strengthens the robustness of evidence of transitions. The transition is identified as Wigner Cat Phases, because of the shape of empirical spectral densities, which depens on the tuneable parameter. These simulated phases in mGOE appear to be an ideal tool to study Eigenstate Thermalization Hypothesis (ETH) and its related transitions, representing a family of physical systems under different localisation and disorder strengths.
keywords:
quantum chaos, spectral fluctuations, random matrix theory, eigenstate thermalization, many-body localisation, quantum thermodynamics
articletype: Paper
1 Introduction
Classical to quantum mechanical correspondence in describing how classically chaotic systems could manifest in quantum mechanics has been an interest with great fascination [berry87, ford92, gutzwiller92, heller93]. In this direction, how to define and quantify quantum chaos is extensively studied [berry77, chirikov88, zurek95, chirikov95, stockmann09, gutzwiller13, haake18]. Among these studies, Bohigas-Giannoni-Schmit (BGS) conjecture [bohigas84] took a prominent place due to its clear practical demonstration of a conceptual foundation that invariant random matrix ensembles of Wigner-Dyson [wishart28, wigner51, wigner55, wigner57, wigner58, wigner67, dyson62a, dyson62b, dyson62c, dyson62d, dyson62e, mehta04, tao, potters20] represents quantum systems whereby their classical counterparts are chaotic in Lyapunov sense and ergodic. This notion is also called quantum chaology due to Sir Berry, but we use the term quantum chaos. In a prior work, Sir Berry’s conjecture[berry77, jarzynski97] made a closely related assertion, that excited eigenfunctions are selected randomly from an ensemble of Gaussian functions.
Researchers become aware of the connections of BGS conjecture and how statistical quantum mechanics could satisfy quantum ergodicity by applying random matrix theory perturbations, known as Eigenstate Thermalization Hypothesis (ETH) [jdeutsch91, srednicki94, srednicki94b, srednicki95, srednicki96, srednicki99], reviewed [dalessio16, deutsch18] and revisited with Loschmidt echo or out of time order correlators (OTOC) measures [foini19, papalardi22, foini25]. The similarity of ETH against BGS is also shown [magan24, weiden25]. In these directions, there has been growing interest for quantum chaos in quantum information theory and its applications [cotler17, magan18, ali20, altland24, borner24, anand24, googleq25].
A localisation phenomenon whereby thermalization wouldn’t be achieved fully or partially is an open area of research, i.e., Many-Body Localisation (MBL) [sierant25] and comprehensive references there-in. The primary open questions are: how MBL phases occur? What are their statistical and physical properties from random matrix perspective? Aiming at this, there are works exploring the approach to ETH via building models and methods for studying these transitions: the transition from integrability to nonintegrability in quantum many-body systems [santos04, chavda14, sierant19, pai20, sierant25].
Figure 1: Spectral densities are numerically identified for different tuning parameters, at (1a) $\mu=0.54$, (1b) $\mu=0.70$, (1c) $\mu=0.88$ and (1d) $\mu=0.98$. These are so-called Wigner Cat Phases due to their M-shaped densities deviating from Wigner’s semi-circle law. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that semicircle law is recovered at small mixtures, i.e., higher $\mu$ values.
We presented a new kind of invariant random matrix ensemble, a mixed version of Wigner-Dyson ensembles. Construction of such ensemble is discussed and along with a principled way of building spectral densities in these mixed ensembles. One of this ensemble, mixed Gaussian Orthogonal Ensemble (mGOE) [suzen20, suzen21] is used in simulating transition to ETH and phases of MBL. Simulations give rise to so-called Wigner Cat Phases due to M-shaped empirical spectral densities. Evolution of the nearest neighbor spacing distribution (NNSD) with unfolding with truncation of spectral outliers [ash14, abu18] and the ratio of two consecutive level spacings (adjcency gap) over a continuous transition that is possible with mGOE are numerically identified. These transitions represent different localisation strengths and MBL phases. Adjcency gaps are preferred way of analysis due to their independence to spectral unfolding procedure [oganesyan07, chavda14]. Sampling from mGOE can be made very finely due to continuous mixture parameter, providing unprecedented resolution to study MBL phases.
In Section 2, we shortly provide a glance of the quantum thermalization and localisation problem. The following Section 3 presents our family of quantum systems represented by invariant mixed Wigner-Dyson ensembles. So-called Wigner Cat Phases are demonstrated with numerical experiments with spectral measures of nearest-neighbour and adjacent ratio level spacings in Section 4. We finally conclude our study in Section 5.
Figure 2: Nearest-neighbour spacing for different $\mu$ tuning are shown, (2a) $\mu=0.54$, (2b) $\mu=0.70$, (2c) $\mu=0.86$, and (2d) $\mu=0.98$. Deviation from Wigner-Dyson distribution at smaller $\mu$ values are demonstrated with lower values indicating heavy-tailed distribution. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that Wigner-Dyson distribution is recovered at small mixtures, i.e., higher $\mu$ values.
2 Quantum thermalization transition to localisation
A definite quantum state $|\Psi(t)\rangle$ can evolve over time given its eigenstates $|\alpha\rangle$ and eigenenergies $E_{\alpha}$ in an exponential fashion with a corresponding observable linear operator $A$,
$|\Psi(t)\rangle=\sum_{\alpha}C_{\alpha}\exp(-iE_{\alpha}t)|\alpha\rangle$
where $\sum_{\alpha}|C_{\alpha}|^{2}=1$, a normalised state. How this system achieves thermal equilibrium, observable reaches to an equilibrium value given temperature $T_{\alpha}$ is shown [jdeutsch91, srednicki94, srednicki94b, srednicki95, srednicki96, srednicki99] by two conditions, so-called Eigenstate Thermalization Hypothesis (ETH):
  1. 1.
    Expectation value $\langle\alpha|A|\alpha\rangle$ is a smooth function of eigenenergies $\Phi(E_{\alpha})$ implying individual eigenstates possess thermal properties, $\langle A\rangle_{T_{\alpha}}=\Phi(E_{\alpha})$, independent of initial values.
  2. 2.
    Given transition from $\alpha$ to $\beta$, diagonal elements of $\langle\alpha|A|\beta\rangle$ should be much larger than off-diagonal elements. This is important as to recover thermal properties for individual eigenstates.
This is also related to quantum ergodicity because ensemble average $\langle A\rangle$ is equal to time average of $A$ if thermalization is achieved. This outstanding and insightful description of quantum thermalization due to Deutsch-Srednicki provided a basis for showing that, quantum system whereby their classical counterparts are chaotic in Lyapunov sense satisfies these two conditions of ETH.
Figure 3: Density of adjacent gap ratios with mean values marked at different degree of mixtures, (3a) $\mu=0.98$, (3b) $\mu=0.92$, (3c) $\mu=0.86$ and (3d) $\mu=0.70$. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that tail is changing at higher mixtures, i.e., lower $\mu$ values.
If we invoke BGS conjecture, as it is shown recently that BGS has better statistical predictive power due its statistical nature [magan24, weiden25]. Hence, Dyson-Wigner matrices in Gaussian Orthogonal Ensemble (GOE) satisfies the termalisation hypothesis, and also descibes quantum chaotic systems. We concentrate on this aspect, that, quantum chaotic systems can thermalize and if we deviate from this behaviour our system will not reach thermal equilibrium. This phenomenon is called Many-Body Localisation (MBL). A more in-depth theoretical review is presented recently by Sierant-Lewenstein et al. [sierant25] and references therein.
The core problem reduced to the statistical identification of deviation from quantum chaos, under which conditions and systems this would occur. A prominent measure in this direction was introduced by Oganesyan-Huse in a seminal work [oganesyan07], adjacent gap ratios on the spectra in investigating the MBL phenomenon. Given $N$ spectral spacings $\delta_{i}$, after sorting eigenenergies, mean adjacent gap ratio $r$ is computed as follows,
$r=\frac{1}{N}\sum_{i=2}^{N}\frac{min(\delta_{i},\delta_{i-1})}{max(\delta_{i},\delta_{i-1})}.$
They found that r takes values of $0.3860$ and $0.5295$ from full localisation to full thermalization. This transition points corresponds to Poisson and Wigner-Dyson distributed quantum systems. Our tunable ensemble as a core tool explained in the next section, can simulate this transition from ETH and phases of MBL. Fascinating phenomenon that this transition corresponds to transition from quantum integrability to nonintegrability and also transition from quantum non-ergodicity to ergodicity.
There are some basic differences among the concepts of chaotic, ergodic and integrable systems [sierant25], but here we concentrate on the chaotic quantum dynamics and the localisation phenomenon based on spectral analysis due to BGS.
3 Invariant Mixing Ensembles: Tuning for Localisation
The primary invariant ensemble that represents a quantum chaotic system as as proposed by BGS conjecture is Gaussian Orthogonal Ensemble (GOE). Its numerical construction was clearly demonstrated by the seminal paper of Edelman-Rao [edelman05], among other generalised ensembles. Here, we concentrate on only GOE. A given matrix $G_{1}$ having entries drawn from Gaussian distribution with mean $\mu=0$ and standard deviation $\sigma$. We can sample GOE and obtain a member of ensemble $A^{GOE}(N)$ with the following algebraic equality [edelman05],a $NxN$ square matrix,
$A^{GOE}(N)=\frac{1}{2}(G_{1}(N)+G_{1}^{T}(N))$
, $G_{1}^{T}(N)$ is the transpose matrix. The sampled matrix $A^{GOE}(N)$ has the following properties.
  1. 1.
    Diagonals with the following distribution: $diag(A^{GOE})\sim\mathcal{N}(0,\sigma)$,
  2. 2.
    Offdiagonals with the following distributions: $offdiag(A^{GOE})\sim\mathcal{N}(0,\sigma/2)$.
If we draw $M$ matrices, we call the resulting set $GOE$ of size $M$ and with square matrix sizes of $N$, denoted by $GOE(M,N)$. By doing this we can estimate bootstrapped confidence intervals on all computed properties [efron94, davison97]. This kind of confidence intervals is not usually practiced in RMT studies but in our context it is important for investigations of the mixed ensembles due to nature of numerical generation of mixed ensembles.
A mixed ensemble, we mix the matrix sizes in the given ensemble of size $M$. A mixed $GOE$ denoted by $mGOE$ can be characterized by one more parameter degree of mixture $\mu$, hence $mGOE(M,N,\mu)$. In the limiting case $\mu=1.0$ mixed ensemble is equivalent to its conventional counterpart, such that $GOE(M,N)=mGOE(M,N,\mu=1.0)$.
Numerical construction of $mGOE$ or any mixed ensemble of the corresponding canonical random matrix ensemble follows the following recipe.
  1. 1.
    There will be $M$ different size random matrices drawned from $GOE(n_{i})$.
  2. 2.
    The number of matrices of size $n_{i}$ is denoted by $m_{i}$ with the constrain $M=\sum_{i}m_{i}$.
  3. 3.
    The degree of mixture $\mu$ can be intepreted as success probability in a Binomial process, that satisfied $M=\sum_{i}m_{i}$ and limiting case $\mu=1.0$ exactly.
  4. 4.
    The set of $n_{i}$ then can be identified by $M$ Binomial trials with the following parametrisation $Binomial(\mu,N)$.
  5. 5.
    Then we drawn matrices from conventional GOE, $M$ times $GOE(n_{i})$, whereby size of matrices comes from the Binomial process.
  6. 6.
    In the large $M$ limit, $N\cdot\mu=\frac{1}{M}\sum_{i}n_{i}$ should hold.
For example, in building $mGOE(M,N,\mu)$, a simple possible case for $M=10$ for $N=1000$ at $\mu=0.8$ a possible set of draws from Binomial process can be the following $Binomial(1000,0.8)$, $[809,813,809,797,808,789,795,810,807,797]$ are the sizes $n_{i}$ of matrices we should generate from $GOE$. Mean of the sizes is $800.1$, in the large $M$ limit this should be exactly $800$.
3.1 Spectral Periodicity: Degenarcy and inducing localisation
The above recipe to generate mixed ensemble will generate many different length spectra. It is important to align spectra of different size matrices. For this purpose, we apply periodic boundary conditions on the resulting spectra. The matrix with size $n_{i}$ will produce $n_{i}$ eigenvalues. Periodicity dictates repeated eigenvalues up to the base size $N$ for mGOE, recall the parameters of the ensemble $mGOE(N,M,\mu)$. This leads to degeneracy in the energy levels.
In investigating localisation phenomenon, MBL, intuitively the degree of mixture $\mu$ plays a role of tunable parameter that drives the representative quantum system of $mGOE$ to localisation away from thermalization at $\mu=1.0$ due to expressed degeneracy and randomness in generating the mixed ensemble via Binomial draws. Effect of such directed randomness is also observed in deep learning [suzen25g].
The theoretical justification of this behaviour lies in mismatched spectral analysis depending on the matrix sizes. This can be imagined via average behaviour of different spectral distribution. Hence, this generates a defect in spectral density in sampling $mGOE$, driving the system to localisation with decreasing $\mu$. From a physical stand point, equivalently this is a kind of inducing a disorder in generating mixed ensemble, amplified by the spectral periodicity and variation of sizes in mixture. This is not a finite-size effect, rather conjectured to be an inherent property of the mixed ensemble.
3.2 Self-consistent spectral unfolding
Unfolding the spectra aimed at removing local fluctuations and making them locally flat on average. It is shown in the literature that computation of spectral nearest-neighbor spacings are sensitive to eigenvalue outliers [abu18]. For this reason, we used interquartile range of eigenvalues without any loss of distributional quality, i.e., perfect agreement with the theoretical distributions in the full thermalization GOE case.
A self-consistent procedure of using multiple different polynomial of a given degree and select the degree such that resulting mean fluctuations are closest to one. Even though adjacent gap ratio is insensitive to unfolding, we used the folded spectrum for all analysis.
Figure 4: Average (Mean) gap ratios over different mixture $\mu$ strengths. These values inform us how localisation changes. Horizontal lines describe reaching to full ETH and full integrability respectively.
4 Numerical Experiments
Numerical investigation of how tuning mGOE gives rise to a deviation from limiting value $\mu=1.0$. At this limiting value ETH is valid in the corresponding representative quantum system’s dynamics. Decreasing $\mu$ parameter represents a more localized states, i.e., MBL. We use matrix order of $N=1000,500$ with ensemble sample size of $M=100$ for quantifying the uncertainties.
The generated dataset using the formulation of mGOE along different $\mu$ at fixed ensemble size is used representing a quantum systems transitioning to MBL. In supporting the transition, we analyze how spectral, nearest-neighbour spacings and adjacent gap ratio densities evolve as a function of the tuning parameter. Additionally, we track mean adjacent gap ratio over tuned parameter to see under which tuning the computed values matches with full ETH and non-chaotic dynamics.
4.1 Wigner Cat Phases: Spectral Densities
Spectral densities are numerically identified for different tuning parameters, in Figures 1a-1d for $\mu=0.54,0.70,0.88,0.98$ respectively. These are so-called Wigner Cat Phases due to their M-shaped densities deviating from Wigner’s semi-circle law. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error-bars, some might be asymmetric as there is no distributional assumption on the errors. We see that semicircle law is recovered at small mixtures, i.e., higher $\mu$ values.
As decreasing $\mu$ represents more localisation, Wigner Cat Phases become more dominant, ”cat ears” approach closer. We identify that this phenomenon originates from the combination of eigenstate degeneracy and randomness creating an effect of a defect, as discussed in the formulation of mGOE generation.
4.2 Spectral Nearest-Neigbour Spacings
Nearest-neighbour spacing for different $\mu$ tuning are shown in Figures 2a-2d for $\mu=0.54,0.70,0.86,0.98$ respectively. Deviation from Wigner-Dyson distribution is demonstrated at smaller $\mu$ values, with lower values showing heavy-tailed distribution. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that Wigner-Dyson distribution is recovered at small mixtures, i.e., higher $\mu$ values. Results supports the notion that MBL phases with heavy-tailed nearest-neighbour spacings are possible.
4.3 Transitions to quantum chaos: Adjecent Gap Ratios
We computed both mean values and distribution of adjacent gap ratios due to Oganesyan-Huse. An advantage of this measure is that, it doesn’t require unfolding the spectra as it isn’t sensitive to the unfolding procedure.
Density of adjacent gap ratios with mean values marked at different degree of mixtures are shown in Figure 3a-3d for $\mu=0.98,0.92,0.86,0.70$ respectively. Uncertainties are computed as usual using bootstrapped $95\%$ confidence intervals.
Mean values of gap ratios inform us about the localisation of the quantum dynamics. Variation of degree of mixture $\mu$ describes the transition from quantum chaos to localisation. mGOE’s can generate such transition in a linear fashion, this is quantified in Figure 4.
We see that in the localized limit, whereby classical counterpart normally behaves like an integrable system, in this case, mGOE at this limit still shows a non-integrable dynamics, see Figure 3a. This observed phenomenon implies a new MBL phase that differs from the interpretation of the established measure of gap ratio.
5 Conclusion
Pedagogically easy to understand ensemble of random matrices is introduced for simulation of localization transition to quantum chaos. Contributions of such mixed ensemble are three fold: (1) advances the notion of matrix ensemble to arbitrarily shaped observable matrices, such as complex networks with varying size sub-components or quantum system as in this work, (2) provides physically plausable toy model that is easy to simulate the transition to quantum chaos with an arbitrarily precise fine tunable parameter for pedagogical and research purposes, and (3) we demonstrated that MBL phases with heavy-tailed nearest-neighbour spacings are possible that still yields mean adjacency-gap value of an integrable system, our results providing spectral data albeit simulated for a new MBL phase, i.e., heavy-tailed localized.
[…]
We are grateful to Y. Süzen for her kind support for the quantum dynamics project. Especially for her encouragement in releasing the computational works as free-software in a Python package for educational use, as well as her recommendations on pedagogical presentation.
[…]
Software toolkit [suzen25ley] and dataset is available.